0Workshop supervisor

F. Frans is the workshop supervisor of NBW. She is paid a fixed salary monthly of N$10 000. To keep her motivated she is entitled to an annual bonus of N$9000. NBV contributes N$9000 to her pension fund annually. She works 9 hours per day from Monday to Friday every week. She is also entitled to 28 days paid vacation leave. There are 12 public holidays in the year for which 9 fall on weekends. Assume a 365 days year.

REQUIRED:

Compute Frans s hourly recovery tariff per hour


1
Expert's answer
2022-03-18T05:11:03-0400

1day = 24hours

30days = x

x = 24*30 = 720 hours.

The bonus he gets annually can be divided by 12 to get monthly bonus as follows:

9000/12 = 750

Every month he bags = 9000+750 = 9750

720 hours = 9750

1hour = x

720x = 9750

x = 9750/720 = N$13.54